OpenStudy (anonymous):

The population of a local species of dragonfly can be found using an infinite geometric series where a1 = 36 and the common ratio is one half. Write the sum in sigma notation and calculate the sum that will be the upper limit of this population.

OpenStudy (anonymous):

@jdoe0001 @whpalmer4 help?

OpenStudy (jdoe0001):

\(\textit{sum of infinite geometric serie}=\Large \Sigma_{i=0}^{\infty}\ a_1\cdot r^i\implies \cfrac{a_1}{1-r}\)

OpenStudy (jdoe0001):

you're given \(\bf a_1\ and \ r\) so plug them in

OpenStudy (anonymous):

so, 36 / 1/2 = 18, so 18 is the sum?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

great, thanks for thr help!:D

OpenStudy (jdoe0001):

yw